# On the separation Łojasiewicz exponent of real semi-analytic sets in two dimension spaces

### 16 Mar 2026 v1

@@ Abstract.
The main aim of the paper is to give the exact formula for the so called separation Łojasiewicz exponent for two real analytic set germs via Newton–Puiseux expansions in two variables. Moreover, we present an effective exponent for the case of two real algebraic sets via their degrees.
@@ 1. Introduction
The Łojasewicz inequalities impressively appeared in 50s years of last century to solve the question of L. Schwartz for distribution division [13, 20, 21]. They have many relations and applications to other branches of mathematics, such as research on local singularities of analytic functions [26], the proof of Thom’s Gradient Conjecture [19], study of infinitely dimensional version in partial differential equations [3], applications in polynomial and tame optimization [11, 12], …
@@
  • There exists $C,r>0$ and $\beta\geq 1$ such that
@@
  • The infimum of such exponent $\beta$ is called separation Łojasiewicz exponent of semi-analytic subsets $A$ and $B$ (at the origin), it is denoted by $\mathscr{L}(A,B)$.
Note that, the separation Łojasiewicz exponent $\mathscr{L}(A,B)$ is a rational number [1, Corollary 2]. The Łojasiewicz exponent represents and compares the growths of functions on two sides of the inequalities [21].
@@ 2.1. Newton polygon relative to an arc
We want to find a Newton-Puiseux root of $f=0$, i.e. a Puiseux series $x=\gamma(y)$ such that $f(\gamma(y),y)=0$. The algorithm has the following steps:
Algorithm 2.1 (Sliding technique).
@@
Note that in the iterative process, the lowest dots on $X=0$ of $\mathbb{P}(f,\phi_{k})$ will move upward. The above algorithm is based on the following lemma.
@@
    If $\tan\theta_{E_{1}}<\rho$ or $\tan\theta_{E_{1}}=\rho$ and $\mathcal{E}_{E_{1}}(c)\neq 0$ then $\mathbb{P}(f,\phi)=\mathbb{P}(f,\psi),$ and therefore $\mathrm{ord}f(\phi(y),y)=\mathrm{ord}f(\psi(y),y).$
@@
If $f=0$ does not have any real Newton–Puiseux root, then $V_{f}=(0,0)$. Take $\epsilon>0$ small sufficiently such that $|y|<\epsilon$, we have
@@
Therefore $d((\phi(y),y),V)\simeq|y|^{\mathrm{ord}d(\phi,V)}$ , for all $|y|<\epsilon$.
On combining this with (2), we deduce the proof of the lemma. ∎
@@ 2.2. Approximations of Puiseux series
Let $f\colon(\mathbb{R}^{2},0)\to(\mathbb{R},0)$ be a real analytic function germ, which is mini-regular in $x$ of order $m$. Suppose that $\phi$ is not Newton–Puiseux root of $f=0$ and $\gamma$ is a final result of sliding of $\phi$ along $f$. Then,
@@
where $\mathcal{E}_{H_{M}}(c)=0$ and $M(X,Y)=f(X+\phi(Y),Y)$. If $c\notin\mathbb{R}$, then by definition of the real approximation (3), the real approximation $\gamma^{\mathbb{R}}$ of series $\gamma$ also has the form
@@
where $g\in\mathbb{R}$ is generic constant. For $f\in\mathbb{K}\{x,y\}$ which is regular in $x$, let $\mathcal{V}_{\mathbb{R}}(f)$ be the set of all real approximations of all Newton-Puiseux roots of $f$.
@@
Let $\phi(y)$ be a non-real Puiseux series. Then the real approximation $\phi^{\mathbb{R}}(y)$ of $\phi(y)$ is the $\rho$-approximation of $\phi$, where $\rho$ is the smallest exponent occurring in $\phi$ with non-real coefficient. It is clear that if $\phi$ is real then the real approximation of $\phi$ is itself.
@@ 3. Separation Łojasewicz exponent of two real analytic subsets
$d(x,V_{f})+d(x,V_{g})\geq Cd(x,V_{f}\cap V_{g})^{\beta},\text{for all}\ \|x\|\leq r.$ (5)
The infimum of the exponents $\beta$ such that Inequality (5) holds, is called the separation Łojasiewicz exponent $\mathscr{L}(V_{f},V_{g})$.
@@
Using the Curve Selection Lemma (see [22]), we can easily prove the following equality
@@
This implies that $\ell(\phi)=1$, which is a contradiction. So, the theorem is proved. ∎
@@ 4. Effective estimation for separation Łojasiewicz exponent
If $V_{f}\subset V_{g}$, then $\mathscr{L}(V_{f},V_{g})=1$. Asume that $V_{f}\not\subset V_{g}$. Applying the classical Łojasiewicz inequality for $F$, there exists $\alpha,c>0$ such that

### 17 Mar 2026 v2

# On the separation Łojasiewicz exponents of real analytic sets in the real plane

@@ Abstract.
The main aim of the paper is to give a formula for computing the separation Łojasiewicz exponents for two real analytic set germs via the Newton–Puiseux expansions of their defining functions. Moreover, we present an effective exponent for the case of two real algebraic sets in terms of their degrees.
@@ 1. Introduction
The Łojasewicz inequalities impressively appeared in the 50s years of last century to solve the question of L. Schwartz for distribution division [13, 20, 21]. They have many relations and applications to other branches of mathematics, such as research on local singularities of analytic functions [26], the proof of Thom’s Gradient Conjecture [19], study of infinitely dimensional version in partial differential equations [3], applications in polynomial and tame optimization [11, 12], …
@@
  • There exist $C,r>0$ and $\beta\geq 1$ such that
@@
  • The infimum of such exponents $\beta$ is called separation Łojasiewicz exponent of semi-analytic subsets $A$ and $B$ (at the origin), it is denoted by $\mathscr{L}(A,B)$.
Note that the separation Łojasiewicz exponent $\mathscr{L}(A,B)$ is a rational number [1, Corollary 2]. The Łojasiewicz exponent represents and compares the growths of functions on two sides of the inequalities [21].
@@ 2.1. Newton polygon relative to an arc
Algorithm 2.1 (The sliding).
@@
The series $\phi_{\infty}$ is called a final result of the sliding of $\phi$ along $f$. The above algorithm is based on the following lemma.
@@
    If $\tan\theta_{E_{1}}<\rho$ or $\tan\theta_{E_{1}}=\rho$ and $\mathcal{E}_{E_{1}}(c)\neq 0$ then $\mathbb{P}(f,\phi)=\mathbb{P}(f,\psi),$ and, therefore $\mathrm{ord}f(\phi(y),y)=\mathrm{ord}f(\psi(y),y).$
@@
If $f=0$ does not have any real Newton–Puiseux root, then $V_{f}=(0,0)$. Taking $\epsilon>0$ small sufficiently such that $|y|<\epsilon$, we have
@@
Therefore, $d((\phi(y),y),V)\simeq|y|^{\mathrm{ord}d(\phi,V)}$ for all $|y|<\epsilon$. Combining this with (2), we deduce the proof of the lemma. ∎
@@ 2.2. Approximations of Puiseux series
Let $f\colon(\mathbb{R}^{2},0)\to(\mathbb{R},0)$ be a real analytic function germ, which is mini-regular in $x$ of order $m$. Suppose that $\phi$ is not a Newton–Puiseux root of $f=0$ and $\gamma$ is a final result of the sliding of $\phi$ along $f$. Then
@@
where $\mathcal{E}_{H_{M}}(c)=0$ and $M(X,Y)=f(X+\phi(Y),Y)$. If $c\notin\mathbb{R}$, then by the definition of the real approximation (3), the real approximation $\gamma^{\mathbb{R}}$ of the series $\gamma$ also has the form
@@
where $g\in\mathbb{R}$ is generic. For $f\in\mathbb{K}\{x,y\}$ which is regular in $x$, let $\mathcal{V}_{\mathbb{R}}(f)$ be the set of all real approximations of all the Newton-Puiseux roots of $f$.
@@
Let $\phi(y)$ be a non-real Puiseux series. Then the real approximation $\phi^{\mathbb{R}}(y)$ of $\phi(y)$ is the $\rho$-approximation of $\phi$, where $\rho$ is the smallest exponent occurring in $\phi$ with non-real coefficient.
@@ 3. Separation Łojasewicz exponent of two real analytic subsets
$d(x,V_{f})+d(x,V_{g})\geq Cd(x,V_{f}\cap V_{g})^{\beta}\text{ for all}\ \|x\|\leq r.$ (5)
@@
Let $\mathscr{L}(V_{f},V_{g})$ denote the separation Łojasiewicz exponent of $V_{f}$ and $V_{g}$. Using the Curve Selection Lemma (see [22]), we can easily prove the following equality
@@
This implies $\ell(\phi)=1$, which is a contradiction. So, the theorem is proved. ∎
@@ 4. Effective estimation for separation Łojasiewicz exponent
If $V_{f}\subset V_{g}$, then $\mathscr{L}(V_{f},V_{g})=1$. Assume that $V_{f}\not\subset V_{g}$. Applying the classical Łojasiewicz inequality to $F$, there exists $\alpha,c>0$ such that